OpenStudy (anonymous):

FAN AND MEDAL (°ロ°)☝ The table shows the results of rolling a number cube, with sides labeled 1 through 6, several times. What is the experimental probability of rolling a 3 or a 6? Express your answer in simplest form. Outcome 1 2 3 4 5 6 Number of times outcome occurred 10 6 4 8 6 6

OpenStudy (anonymous):

it's 2/5

OpenStudy (anonymous):

how do u know

OpenStudy (anonymous):

or how did u find it

OpenStudy (anonymous):

@wontonsilver

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

give me a sec to explain

OpenStudy (anonymous):

10+6/40=16/40=8/20=4/10

OpenStudy (anonymous):

then the answer is 2/5 simplfied

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

can i get a medal

OpenStudy (anonymous):

and fan me plox